OpenStudy (anonymous):

if cos feta=-3/5 and pi

OpenStudy (anonymous):

you need the "half angle" formula for cosine. do you know it?

OpenStudy (anonymous):

it is \[cos(\frac{\theta)}{2}=\pm\sqrt{\frac{1+cos(\theta)}2}\]

OpenStudy (anonymous):

so just replace theta by \[\frac{-3}{5}\]

OpenStudy (anonymous):

you know by the condition \[\pi<\theta<\frac{3\pi}{2}\] that you are in quadrant II so cosine is negative there. make sure you take the negative square root, not the positive.

OpenStudy (anonymous):

excuse me i did not mean replace theta by -3/5 i meant replacse cos(t) by -3/5

OpenStudy (anonymous):

you get \[\sqrt{\frac{1-\frac{3}{5}}{2}}\]

OpenStudy (anonymous):

=\[\sqrt{\frac{\frac{2}{5}}{2}}\]

OpenStudy (anonymous):

=\[-\sqrt{\frac{2}{10}}\]

OpenStudy (anonymous):

=\[\sqrt{\frac{1}{5}}\] if my arithmetic is right